OpenStudy (anonymous):

Write an equation for a line perpendicular to y = -2 x -3 and passing through the point (-6,-8)

OpenStudy (anonymous):

y = 1/2x-5

OpenStudy (mathmale):

Find the slope of y = -2x-3. Hint: You can do this by inspection. No algebra necessary! Find the negative reciprocal of that slope.

OpenStudy (anonymous):

\[y=\frac{ 1 }{ 2 }x - 5\]

OpenStudy (anonymous):

I remember how to get the slope, but I don't remember what to do with the 2 points? how did u get -5?

OpenStudy (mathmale):

Using this new slope, write the equation of the line through (-6,-8). @trueseminole : New to OpenStudy? If so, welcome. Note, however, that one of the rules of this site is that we do NOT do the work of others for them and then hand out the answer without the involvement of the person asking the question. Please help the other person find his or her own answer. Thank you.

OpenStudy (mathmale):

Jessica: what is the slope of the given line? Type it here, please.

OpenStudy (anonymous):

OH, sorry..yes the slope is 1/2

OpenStudy (mathmale):

The slope of the given line is actually m = -2. What is the "negative reciprocal" of that? If you're not sure what "negative reciprocal" means, please ask.

OpenStudy (anonymous):

right, so the negative reciprocal is 1/2

OpenStudy (anonymous):

Oh I see what you meant... The slope is -2...I jumped straight to the neg. recip.

OpenStudy (mathmale):

Right. Now, we know that this new line, the one perpendicular to the given line, passes through the point (-6,-8). Use the point-slope form of the equation for a straight line to determine the equation of the new line. Before you do that, verify that you know what the "point-slope" form looks like.

OpenStudy (anonymous):

point slope? y-y1=m(x-x1)... I'm not sure, but I think that's what it is?

OpenStudy (mathmale):

Yes. This is an extremely important formula, and so it's worth recording, studying and using it. You know the slope, m, is 1/2, and that the point in question is (-6,-8). Please plug this data into the equation you've just typed.

OpenStudy (anonymous):

So y-(-8)=1/2(x-(-6))

OpenStudy (mathmale):

If only you'd put parentheses around that 1/2, your result would be perfect. You could, if you wanted, go furtehr forward by simplifying what you've typed in. but you've shown that you know what to do to arrive at this point. Nice work, and thanks for your patience and persistence.

OpenStudy (anonymous):

ok, thank you so much for your help... I believe I know what do from here :)

OpenStudy (mathmale):

My pleasure. All the best to you! Bye!